# Communication-Efficient Decentralized Optimization via Double-Communication Symmetric ADMM

### v1

@@ Appendix C Proof of Theorem 2
The proof is based on the notations of global form (5). First, we denote vectors
$z=\begin{pmatrix}u\\ v\end{pmatrix},\quad w=\begin{pmatrix}u\\ v\\ \lambda\end{pmatrix},\quad F(w)=\begin{pmatrix}-A^{\top}\lambda\\ B^{\top}\lambda\\ Au-Bv\end{pmatrix}$ (16)
@@
$\widetilde{z}^{(t)}=\begin{pmatrix}\widetilde{u}^{(t)}\\ \widetilde{v}^{(t)}\end{pmatrix},\quad\widetilde{w}^{(t)}=\begin{pmatrix}\widetilde{u}^{(t)}\\ \widetilde{v}^{(t)}\\ \widetilde{\lambda}^{(t)},\end{pmatrix},\quad h(z)=f(x)+g(y)$ (18)
@@
$w^{(t+1)}=w^{(t)}-M(w^{(t)}-\widetilde{w}^{(t)}).$ (21)
@@
$\lambda_{\max}(H)\leq\theta,\quad\lambda_{\min}(G)\geq 2(1-r)\rho$ (23)
@@
combining above, we get $\lambda_{\max}(H)\leq\theta$. Also, we have
@@
This equivalent form of solution set $\Omega^{*}$ follows from [14] by the definition of subdifferentials:
$\Omega^{*}=\bigcap_{w\in\mathbb{R}^{4nd}}\left\{\widehat{w}\,\middle|\,h(z)-h(\widehat{z})+\langle w-\widehat{w},F(w)\rangle\geq 0\right\}.$ (29)
Then we give the following important lemma, whose proof follows directly from the proof of Theorem 2 and Theorem 3 in [3] and the above form of solution set:
@@
$\left\|w^{(t+1)}-w^{*}\right\|_{H}^{2}\leq\left\|w^{(t)}-w^{*}\right\|_{H}^{2}-\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2},\quad\forall w^{*}\in\Omega^{*},$ (30)
@@
$T_{\text{KKT}}(w):=\begin{pmatrix}\partial f(u)-A^{\top}\lambda\\ \partial g(v)+B^{\top}\lambda\\ Au-Bv\end{pmatrix}$. (31)
@@
The sequences $\{w^{(t)}\}$ and $\{\widetilde{w}^{(t)}\}$ satisfy
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(w^{(t+1)}))\leq\frac{\delta}{\theta}\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2},$
@@
By Proposition 2.1 in [14] and our notations, we derive the following inequality characterizing the subproblems:
@@
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(w))=\mathrm{dist}^{2}(0,T_{1}(w))+\mathrm{dist}^{2}(0,T_{2}(w))+\mathrm{dist}^{2}(0,T_{3}(w))$ (34)
where $T_{1}(w)=\partial f(u)-A^{\top}\lambda$, $T_{2}(w)=\partial g(v)+B^{\top}\lambda$, $T_{3}(w)=Au-Bv$.
@@
$\mathrm{dist}(0,T_{1}(w^{(t+1)}))$ $\leq\left\|A^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t+1)})-Q(\widetilde{u}^{(t)}-u^{(t)})\right\|$ (35)
@@
$\mathrm{dist}(0,T_{2}(w^{(t+1)}))$ $\leq\left\|B^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t+1)})-(Q+\beta B^{\top}B)(\widetilde{v}^{(t)}-v^{(t)})+rB^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t)})\right\|$ (37)
@@
$\mathrm{dist}(0,T_{3}(w^{(t+1)}))$ $=\left\|\frac{1}{\beta}(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-B(v^{(t)}-\widetilde{v}^{(t)})\right\|.$ (38)
Substituting  (35) (37) (38) into (34) while using the definition of $A$, $B$, $Q$, we get
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(w^{(t+1)}))$ $\leq\left\|A^{\top}\left[r(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-\beta B(v^{(t)}-\widetilde{v}^{(t)})\right]-Q(\widetilde{u}^{(t)}-u^{(t)})\right\|^{2}$ (39)
@@
$\leq\delta\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|^{2}\ \leq\frac{\delta}{\phi}\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2}.$
@@
Because $\Omega^{*}$ is a closed convex set, there exists a $w^{*}_{t}\in\Omega^{*}$ satisfying
$\operatorname{dist}_{H}(w^{(t)},\Omega^{*})=\left\|w^{(t)}-w^{*}_{t}\right\|_{H}.$ (40)
@@
$\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}\geq$ $\sqrt{\frac{\phi}{\delta}}\mathrm{dist}(0,T_{\text{KKT}}(w^{(t+1)}))$ (41)
$\geq$ $\sqrt{\frac{\phi}{c^{2}\delta}}\operatorname{dist}(w^{(t+1)},\Omega^{*})$
$\geq$ $\sqrt{\frac{\phi}{c^{2}\delta\theta}}\operatorname{dist}_{H}(w^{(t+1)},\Omega^{*}).$
@@
$(1+\epsilon)\operatorname{dist}_{H}^{2}(w^{(t+1)},\Omega^{*})\leq$ $\left\|w^{(t+1)}-w^{*}_{t}\right\|_{H}^{2}+\epsilon\operatorname{dist}_{H}^{2}(w^{(t+1)},\Omega^{*})$ (42)
$\leq$ $\left\|w^{(t+1)}-w^{*}_{t}\right\|_{H}^{2}+\left\|w^{(t)}-\widetilde{w}^{(t)}\right\|_{G}^{2}$
$\leq$ $\left\|w^{(t)}-w^{*}_{t}\right\|_{H}^{2}=\operatorname{dist}_{H}^{2}(w^{(t)},\Omega^{*})$
@@
To prove the R-linear rate of suboptimality, first, from the same approach of proving Corollary 2.1 in [2] we are able to prove the R-linear rate of $\left\|w^{(t)}-w^{\infty}\right\|$. And a simple corollary from Lemma 2 is that $\left\|w^{(t)}-\tilde{w}^{(t)}\right\|$ and $\left\|Au^{(t+1)}-Bv^{(t+1)}\right\|$ also converges R-linearly.
Then we substitute $u^{\infty}$ and $v^{\infty}$ into (32) and  (33):
@@
the above inequalities and the R-linear convergence of $\left\|w^{(t)}-w^{\infty}\right\|$, $\left\|w^{(t)}-\tilde{w}^{(t)}\right\|$, $\left\|Au^{(t+1)}-Bv^{(t+1)}\right\|$ lead to the R-linear convergence of suboptimality, which finishes the proof.
Appendix D Additional Experimental Results
In this section, we report experimental results on a 30-agent sparsely connected random graph with edge probability $p=0.2$. As expected, all algorithms perform worse on this sparsely connected topology, consistent with the theoretical results. Nevertheless, our proposed DS-ADMM still demonstrates the best overall performance.
D.1 Lasso Regression
Figure 3: Performance on the Lasso regression task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
D.2 SVM Classification
Figure 4: Performance on the SVM classification task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.

### v2

@@ Appendix C Proof of Theorem 2
The proof is based on the notations of global form equation 7. First, we denote vectors
$z=\begin{pmatrix}u\\ v\end{pmatrix},\quad\theta=\begin{pmatrix}u\\ v\\ \lambda\end{pmatrix},\quad F(\theta)=\begin{pmatrix}-A^{\top}\lambda\\ B^{\top}\lambda\\ Au-Bv\end{pmatrix}$ (20)
@@
$\widetilde{z}^{(t)}=\begin{pmatrix}\widetilde{u}^{(t)}\\ \widetilde{v}^{(t)}\end{pmatrix},\quad\widetilde{\theta}^{(t)}=\begin{pmatrix}\widetilde{u}^{(t)}\\ \widetilde{v}^{(t)}\\ \widetilde{\lambda}^{(t)},\end{pmatrix},\quad h(z)=f(x)+g(y)$ (22)
@@
$\theta^{(t+1)}=\theta^{(t)}-M(\theta^{(t)}-\widetilde{\theta}^{(t)}).$ (25)
@@
$\lambda_{\max}(H)\leq\zeta,\quad\lambda_{\min}(G)\geq 2(1-r)\rho$ (27)
@@
combining above, we get $\lambda_{\max}(H)\leq\zeta$. Also, we have
@@
This equivalent form of solution set $\Omega^{*}$ follows from (He et al., 2016) by the definition of subdifferentials:
$\Omega^{*}=\bigcap_{\theta\in\mathbb{R}^{4nd}}\left\{\widehat{\theta}\,\middle|\,h(z)-h(\widehat{z})+\langle\theta-\widehat{\theta},F(\theta)\rangle\geq 0\right\}.$ (33)
Then we give the following important lemma, whose proof follows directly from the proof of Theorem 2 and Theorem 3 in (Bai et al., 2018) and the above form of solution set:
@@
$\left\|\theta^{(t+1)}-\theta^{*}\right\|_{H}^{2}\leq\left\|\theta^{(t)}-\theta^{*}\right\|_{H}^{2}-\left\|\theta^{(t)}-\widetilde{\theta}^{(t)}\right\|_{G}^{2},\quad\forall\theta^{*}\in\Omega^{*},$ (34)
@@
$T_{\text{KKT}}(\theta):=\begin{pmatrix}\partial f(u)-A^{\top}\lambda\\ \partial g(v)+B^{\top}\lambda\\ Au-Bv\end{pmatrix}$. (35)
@@
The sequences $\{\theta^{(t)}\}$ and $\{\widetilde{\theta}^{(t)}\}$ satisfy
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(\theta^{(t+1)}))\leq\frac{\delta}{\phi}\left\|\theta^{(t)}-\widetilde{\theta}^{(t)}\right\|_{G}^{2},$
@@
By Proposition 2.1 in (He et al., 2016) and our notations, we derive the following inequality characterizing the subproblems:
@@
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(\theta))=\mathrm{dist}^{2}(0,T_{1}(\theta))+\mathrm{dist}^{2}(0,T_{2}(\theta))+\mathrm{dist}^{2}(0,T_{3}(\theta))$ (38)
where $T_{1}(\theta)=\partial f(u)-A^{\top}\lambda$, $T_{2}(\theta)=\partial g(v)+B^{\top}\lambda$, $T_{3}(\theta)=Au-Bv$.
@@
$\mathrm{dist}(0,T_{1}(\theta^{(t+1)}))$ $\leq\left\|A^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t+1)})-Q(\widetilde{u}^{(t)}-u^{(t)})\right\|$ (39)
@@
$\mathrm{dist}(0,T_{2}(\theta^{(t+1)}))$ $\leq\left\|B^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t+1)})-(Q+\beta B^{\top}B)(\widetilde{v}^{(t)}-v^{(t)})+rB^{\top}(\widetilde{\lambda}^{(t)}-\lambda^{(t)})\right\|$ (41)
@@
$\mathrm{dist}(0,T_{3}(\theta^{(t+1)}))$ $=\left\|\frac{1}{\beta}(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-B(v^{(t)}-\widetilde{v}^{(t)})\right\|.$ (42)
Substituting  equation 39 equation 41 equation 42 into equation 38 while using the definition of $A$, $B$, $Q$, we get
$\mathrm{dist}^{2}(0,T_{\text{KKT}}(\theta^{(t+1)}))$ $\leq\left\|A^{\top}\left[r(\lambda^{(t)}-\widetilde{\lambda}^{(t)})-\beta B(v^{(t)}-\widetilde{v}^{(t)})\right]-Q(\widetilde{u}^{(t)}-u^{(t)})\right\|^{2}$ (43)
@@
$\leq\delta\left\|\theta^{(t)}-\widetilde{\theta}^{(t)}\right\|^{2}\ \leq\frac{\delta}{\phi}\left\|\theta^{(t)}-\widetilde{\theta}^{(t)}\right\|_{G}^{2}.$
@@
Because $\Omega^{*}$ is a closed convex set, there exists a $\theta^{*}_{t}\in\Omega^{*}$ satisfying
$\operatorname{dist}_{H}(\theta^{(t)},\Omega^{*})=\left\|\theta^{(t)}-\theta^{*}_{t}\right\|_{H}.$ (44)
@@
$\left\|\theta^{(t)}-\widetilde{\theta}^{(t)}\right\|_{G}\geq$ $\sqrt{\frac{\phi}{\delta}}\mathrm{dist}(0,T_{\text{KKT}}(\theta^{(t+1)}))$ (45)
$\geq$ $\sqrt{\frac{\phi}{c^{2}\delta}}\operatorname{dist}(\theta^{(t+1)},\Omega^{*})$
$\geq$ $\sqrt{\frac{\phi}{c^{2}\delta\zeta}}\operatorname{dist}_{H}(\theta^{(t+1)},\Omega^{*}).$
@@
$(1+\epsilon)\operatorname{dist}_{H}^{2}(\theta^{(t+1)},\Omega^{*})\leq$ $\left\|\theta^{(t+1)}-\theta^{*}_{t}\right\|_{H}^{2}+\epsilon\operatorname{dist}_{H}^{2}(\theta^{(t+1)},\Omega^{*})$ (46)
$\leq$ $\left\|\theta^{(t+1)}-\theta^{*}_{t}\right\|_{H}^{2}+\left\|\theta^{(t)}-\widetilde{\theta}^{(t)}\right\|_{G}^{2}$
$\leq$ $\left\|\theta^{(t)}-\theta^{*}_{t}\right\|_{H}^{2}=\operatorname{dist}_{H}^{2}(\theta^{(t)},\Omega^{*})$
@@
To prove the R-linear rate of suboptimality, first, from the same approach of proving Corollary 2.1 in (Bai et al., 2019) we are able to prove the R-linear rate of $\left\|\theta^{(t)}-\theta^{\infty}\right\|$. And a simple corollary from Lemma 2 is that $\left\|\theta^{(t)}-\tilde{\theta}^{(t)}\right\|$ and $\left\|Au^{(t+1)}-Bv^{(t+1)}\right\|$ also converges R-linearly.
Then we substitute $u^{\infty}$ and $v^{\infty}$ into equation 36 and  equation 37:
@@
the above inequalities and the R-linear convergence of $\left\|\theta^{(t)}-\theta^{\infty}\right\|$, $\left\|\theta^{(t)}-\tilde{\theta}^{(t)}\right\|$, $\left\|Au^{(t+1)}-Bv^{(t+1)}\right\|$ lead to the R-linear convergence of suboptimality, which finishes the proof.
Appendix D Additional Experimental Details
D.1 Parameter Sensitivity Analysis
We summarize the hyperparameters used in the DS-ADMM algorithm and clarify their respective roles.
Proximal parameter $\tau$.
The proximal term ensures that the matrix $Q$ is positive definite, which is required in order to establish linear convergence. Throughout our experiments, we set $\tau=0.01$.
Augmented Lagrangian parameter $\beta$.
The parameter $\beta$ has a major influence on the convergence speed. One may either adopt the adaptive tuning strategy proposed in (Boyd et al., 2011) or evaluate several candidate values and select the one that performs best empirically.
Dual update step sizes $(r,s)$.
The convergence region and admissible ranges for the symmetric dual step sizes have been characterized in prior work (Bai et al., 2018). Following these theoretical guidelines, we fix $s=1$ and $r=0.99$, which consistently yield stable and efficient performance.
D.2 Additional Experimental Results
We report experimental results on a 30-agent sparsely connected random graph with edge probability $p=0.2$. As expected, all algorithms perform worse on this sparse topology, consistent with theoretical predictions. Nevertheless, our proposed DS-ADMM retains the best overall performance.
Figure 3: Lasso performance on the 30-agent sparse random graph ($p=0.2$). Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Figure 4: SVM performance on the 30-agent sparse random graph ($p=0.2$). Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
We further evaluate the Lasso problem on a 100-agent network under two connectivity levels, with edge probabilities $p=0.5$ and $p=0.2$.
Figure 5: Lasso performance on the 100-agent random graph ($p=0.5$). Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Figure 6: Lasso performance on the 100-agent random graph ($p=0.2$). Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Then we test performance when the dataset is randomly partitioned among agents, using $n=30$ and $n=100$ with graph edge probability $p=0.5$.
Figure 7: Lasso performance with random data partition on the 30-agent random graph ($p=0.5$). Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Figure 8: Lasso performance with random data partition on the 100-agent random graph ($p=0.5$). Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Finally, we test the performance on a ring network with $50$ agents.
Figure 9: Lasso performance on the 50-agent ring topology. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Appendix E LLM Usage
We used a large language model (LLM) to help polish the writing and enhance readability. The authors are solely responsible for the technical content, analysis, and conclusions.